# Advancing Compositional Awareness in CLIP with Efficient Fine-Tuning

### v1

@@ D.4 Other methods do not generalize
From the results associated with SugarCrepe++ and SugarCrepe, one sees that other compositionality enhancing methods like SVLC-R, DAC etc., perform relatively well on SugarCrepe but do not generalize to SugarCrepe++. We believe this is due to the way these methods operate. Specifically, most of these methods create hard-negatives aimed at certain attributes like color, position, etc, which is exactly what benchmarks like ARO, and SugarCrepe (ARO addresses the same task, but SugarCrepe generates more fluent and sensical negatives) test for. For SugarCrepe++, in addition to the evaluation in SugarCrepe, an additional lexically different query text is introduced and since these methods do not account for general compositional/structural variations in their hard-negative, they fall short on SugarCrepe++. On the other hand, the random word swapping in CLIC makes our method focus not on the specific attributes but the general composition of different words in the query text. This makes our method generalize better and work equally well for both SugarCrepe++ and SugarCrepe.
@@ Appendix E Additional experiments
As we only show averaged ITT and TOT numbers across different sets of SugarCrepe++ and SugarCrepe in the main paper, the full versions for ViT-B/32 models can be found in Table 8 and Table 9 resp. Similarly, for ViT-B/16 (omitted in the main part) and ViT-L/14, detailed results are shown in Table 16 and Table 17, respectively. Detailed compositionality results for the newer CLIP versions in CLIPA, EVA02-CLIP and CLIPS with their respective better versions achieved via CLIC can be found in Table 18. In Table 15, we give dataset wise accuracy of all the large models for the ZS-10 setting. WinoGround results for large model can be found in Table 19. The pre-trained checkpoints used for fine-tuning with CLIC are listed in Table 14.
Improvements across architectures
@@
In Table 3 in the main part of the paper, we show how CLIC improves compositionality and retrieval for differently pre-trained models. We visualize these improvements in Figure 5. Fine-tuning with CLIC yields improvements on all models for ITT on SugarCrepe++ and SugarCrepe. The improvements in case of CLIP are as high as 7% and as much as 9% for CLIPS.
Ablating CLIC
In Table 10, we analyse the components of CLIC (), where we concatenate two images, along with its analogous version using only a single image (). For the single image version of our CLIC, fine-tuning on the Laion data () improves SugarCrepe results a bit and marginally improves for some cases of SugarCrepe++. Adding a single hard-negative (), drops the downstream performance without any improvement in compositionality. Adding more positives ( and ) improves ITT for SugarCrepe++ but leads to a decay in downstream and SugarCrepe numbers. Note that is identical to in this case, as one only has a single image. Adding the standard CLIP loss in recovers downstream performance back.
The ablations for CLIC reveals that starting from (the standard pre-trained CLIP model) and training with concatenated images and a single caption manages to improve some categories TOT in SugarCrepe++, but does not improve ITT scores on SugarCrepe++. Further adding a single hard-negative () does not consistently improve results on SugarCrepe++ but improves SugarCrepe results, suggesting that the hard-negative distractor might help the model to detect changes, but does not help the model to distinguish between semantically similar sentences to ones with different meanings. Furthermore, integrating the Uni-Modal loss function (6) () either improves or maintains performance in both ITT and TOT tasks, as it helps the model discover syntactic changes that do not change the meaning of the caption. In addition, since the model has been trained only on concatenated images, but standard benchmarks measure single-image tasks, we can see the recovery of the model during the standard CLIP loss iterations (, this is the CLIC-LAION version). Finally, our method demonstrates advantages over the single image-text () baseline (uses the additional positives and and the clip loss iterates).
Ablating training steps
@@
In Table 11, we show how the proposed CLIC changes with more training steps. Here, we use MS-COCO for training which means 5 epochs corresponds to 2.5 Laion epochs of our dataset. From the table, we see increasing gains in SugarCrepe++ numbers when increasing the number of training step. However, this comes with a marginal degradation in zero-shot classification, which goes down from 63.2% to 62.7% on ImageNet as we move from 2 to 5 epochs. The same trend holds for our 1M set of Laion, where training for more than 1 epochs led to decay on ImageNet and ZS-10 numbers with marginal gains on SugarCrepe++. Hence, for our runs in the main part of the paper with Laion (CLIC-LAION) we only train for 1 epoch.
Ablating freezing of model components
@@
In Table 12, we show how the proposed CLIC (frozen vision encoder) yields, on average, the best compositionality without sacrificing downstream performance. From the table, it is clear that freezing just the text encoder leads to the least amount of improvement on compositional benchmarks. Fine-tuning the whole model works better (in terms of retrieval) but the gains on SugarCrepe++ are smaller than freezing the vision encoder and it also degrades zero-shot classification performance. From these results, we infer that the most improving configuration is freezing the vision encoder, which is the final setting for CLIC. This also highlights that substantial gains can be made by improving the text encoder alone.
Ablating number of additional positives
@@
In Table 13, we show the impact of adding additional positives for each concatenated image ( and from Section 3.1). Notably, increasing the number of positives has opposing effects on the TOT and ITT scores. In addition, after two positives, the effects on ITT in SugarCrepe++ become minimal, so we chose to add only two positives.
More retrieval results
@@
In the main part of the paper, all results on MS-COCO and Flickr30k retrieval are reported with Recall@5. In Table 20, we additionally report Recall@1 for all ViT-B/32 models from Table 1. The overall ordering of methods is similar to Recall@5. Similarly, in Table 19, we additionally report Recall@1 for all ViT-L/14 models from Table 3.
@@
Downstream Evaluations | Sugarcrepe++ | Sugarcrepe | ||||||||||
IMNET | ZS-10 | Text | Image | ITT | TOT | ITT | TOT | ITT | ITT | ITT | ||
Single image Baselines (Non-Concat) | ||||||||||||
Ablating CLIC (Concat) | ||||||||||||
+CLIP iterate (CLIC) | 61.7 | 61.0 | 75.9 | 60.0 | 75.6 | 60.1 | 61.1 | 27.9 | 84.5 | 84.0 | 73.7 |

### v2

@@ D.4 Other methods do not generalize
From the results on SugarCrepe++ and SugarCrepe, one sees that other compositionality enhancing methods like SVLC-R, DAC, etc., perform relatively well on SugarCrepe but fail to generalize to SugarCrepe++. We believe this is due to the way these methods operate. Specifically, most of these methods create hard-negatives aimed at certain attributes such as color, position, or size (for example, modifying P1: "a woman plays a black guitar" to N: "a woman plays a green guitar"). This aligns closely with what benchmarks like ARO, and SugarCrepe evaluate (with SugarCrepe generating more fluent and sensical negatives).
While such approaches can indeed improve performance on lexical understanding, they fail to capture word replacements that are not in the dictionary and fall outside of a predefined structure. More importantly, we believe these changes do not foster a deeper, more nuanced understanding of compositionality within the model. True compositional reasoning requires distinguishing between examples that differ not only lexically but also syntactically and semantically, for instance, recognizing that N2: “a woman plays a plastic guitar” is negative, whereas P2: “a guitar is being played by a woman” is positive. These cases involve more complex syntactic and semantic variations and move beyond lexical differences.
For SugarCrepe++, in addition to the evaluation in SugarCrepe, an additional lexically different text is introduced, and since these methods do not account for general compositional/structural variations in their hard-negative, they fall short on SugarCrepe++. On the other hand, the random word swapping in CLIC makes our method focus not on the specific attributes but the general composition of different words in the text. This makes our method generalize better and work equally well for both SugarCrepe++ and SugarCrepe.
D.5 Detailing the LLaVA experiment
As LLaVA uses only the vision encoder from CLIP models trained at a higher image resolution ( pixels), we performed an additional fine-tuning of ViT-L/14 with CLIC, this time also unfreezing the vision encoder. We employed the PixelProse dataset and reduced the effective batch size to , while keeping all other settings identical to those in Section˜D.1.
Subsequently, we fine-tuned LLaVA-1.5 using Vicuna-7b as the LLM. Initially, only the multi-modal projector was fine-tuned, followed by fine-tuning of both the projector and the LLM. This procedure follows the standard setup from the official LLaVA codebase222https://github.com/haotian-liu/LLaVA/tree/main. To ensure a fair comparison, we conducted these experiments using both the original CLIP and our CLIC vision encoders. Hence, the two resulting models in Table˜6 are both fine-tuned by us, making the subsequent comparison fair.
For the evaluation reported in Table˜6, we adopt the VQAScore from Lin et al. [30] for both SugarCrepe++ and WinoGround. For standard VLM benchmarks, we select a subset of tasks from the original LLaVA codebase. Specifically, we assess question answering using accuracy on GQA [19] and TextVQA [47], chain-of-thought reasoning with ScienceQA-Images (SQA-I) [35], and perception via MME [54]. For all evaluations we use the default setup from the original LLaVA codebase.
@@ Appendix E Additional experiments
Extension of results from the main paper As we only show averaged ITT and TOT numbers across different sets of SugarCrepe++ and SugarCrepe in the main paper, the full versions for ViT-B/32 models can be found in Table˜17 and Table˜18 resp. Similarly, for ViT-B/16 (omitted in the main part) and ViT-L/14, detailed results are shown in Table˜24 and Table˜25, respectively. Detailed compositionality results for the newer CLIP versions in CLIPA, EVA02-CLIP and CLIPS with their respective better versions achieved via CLIC can be found in Table˜26. In Table˜23, we give dataset wise accuracy of all the large models for the ZS-10 setting. WinoGround results for large model can be found in Table˜27. The pre-trained checkpoints used for fine-tuning with CLIC are listed in Table˜22.
@@
Improvements across architectures In Table˜3 in the main part of the paper, we show how CLIC improves compositionality and retrieval for differently pre-trained models. We visualize these improvements in Figure˜5. Fine-tuning with CLIC yields improvements on all models for ITT on SugarCrepe++ and SugarCrepe. The improvements in case of CLIP are as high as 7% and as much as 9% for CLIPS.
@@
Ablating training steps In Table˜19, we show how the proposed CLIC changes with more training steps. Here, we use MS-COCO for training which means 5 epochs corresponds to 2.5 Laion epochs of our dataset. From the table, we see increasing gains in SugarCrepe++ numbers when increasing the number of training step. However, this comes with a marginal degradation in zero-shot classification, which goes down from 63.2% to 62.7% on ImageNet as we move from 2 to 5 epochs. The same trend holds for our 1M set of Laion, where training for more than 1 epochs led to decay on ImageNet and ZS-10 numbers with marginal gains on SugarCrepe++. Hence, for our runs in the main part of the paper with Laion (CLIC-LAION) we only train for 1 epoch.
@@
Ablating freezing of model components In Table˜20, we show how the proposed CLIC (frozen vision encoder) yields, on average, the best compositionality without sacrificing downstream performance. From the table, it is clear that freezing just the text encoder leads to the least amount of improvement on compositional benchmarks. Fine-tuning the whole model works better (in terms of retrieval) but the gains on SugarCrepe++ are smaller than freezing the vision encoder and it also degrades zero-shot classification performance. From these results, we infer that the most improving configuration is freezing the vision encoder, which is the final setting for CLIC. This also highlights that substantial gains can be made by improving the text encoder alone.
@@
Ablating number of additional positives In Table˜21, we show the impact of adding additional positives for each concatenated image ( and from Section˜3.1). Notably, increasing the number of positives has opposing effects on the TOT and ITT scores. In addition, after two positives, the effects on ITT in SugarCrepe++ become minimal, so we chose to add only two positives.
Error bars To show the resilience to randomness of CLIC, we conduct three runs for two versions of CLIC presented in Table˜1 and Table˜2 (CLIC-LAION and CLIC-RedCaps). We report the mean and standard deviation in Table˜10 and Table˜11, respectively. The standard deviation is small, particularly when considering that the results in the paper are rounded to one decimal place.
| SugarCrepe++ [11] | WinoGround [52] | SugarCrepe [18] | ||||||||
| Replace | Swap | Text | Image | Group | Add | Replace | Swap | |||
| Method | ITT | TOT | ITT | TOT | Score | Score | Score | ITT | ITT | ITT |
| CLIP | 69.5 | 60.5 | 45.7 | 25.9 | 31.2 | 11.0 | 8.7 | 72.9 | 80.0 | 62.7 |
| CLIC-LAION | 75.34 0.18 | 60.10 0.11 | 61.55 0.38 | 27.86 0.08 | 31.75 0.20 | 11.83 0.31 | 9.33 0.31 | 84.56 0.09 | 83.84 0.14 | 73.79 0.12 |
| CLIC-RedCaps | 76.18 0.16 | 57.82 0.17 | 61.60 0.07 | 23.34 0.19 | 32.00 0.61 | 11.67 0.12 | 10.08 0.24 | 86.27 0.15 | 84.80 0.04 | 72.54 0.04 |
| Classification | ||||||
| Method | IMNET | ZS10 | COCO | F30K | COCO | F30K |
| CLIP | 63.3 | 61.4 | 74.1 | 95.1 | 54.6 | 83.5 |
| CLIC-LAION | 61.72 0.02 | 60.89 0.06 | 75.92 0.17 | 95.10 0.08 | 60.32 0.64 | 86.70 0.03 |
| CLIC-RedCaps | 62.41 0.09 | 60.22 0.03 | 76.00 0.09 | 95.67 0.05 | 59.42 0.04 | 86.27 0.02 |
Long captions retrieval To evaluate the influence of exploiting superficial artifacts such as potential abrupt topic changes, generated by our concatenation scheme, we compared the retrieval performance on longer captions (up to 200 tokens) between the original model, other compositionality-enhancing methods, and our fine-tuned models (CLIC) on the ViT-B/32 architecture. This setup is adapted from long-retrieval as previously done in [58]. We truncate the caption to the nearest punctuation (within context-length) as CLIP has a context length of 77. It is important to note that these longer captions describe a single scene but consist of multiple sentences. This design choice was deliberate: it helps us determine if our CLIC models have inadvertently overfit to abrupt topic changes or even to concatenation with punctuation. As can be seen in Table˜12, only CLIC and TripletCLIP outperform the base model for both image and text retrieval, whereas DAC-LLM and SVLC-R+L are worse.
| Method | ||
| CLIP | 83.2 | 79.5 |
| NegCLIP | 82.6 | 80.7 |
| DAC-LLM | 67.6 | 72.5 |
| SVLC-R+L | 75.0 | 74.8 |
| TripletCLIP | 84.1 | 81.5 |
| CLIC-COCO | 85.9 | 81.4 |
| CLIC-LAION | 83.9 | 79.6 |
| CLIC-RedCaps | 84.7 | 79.9 |
@@
More retrieval results In the main part of the paper, all results on MS-COCO and Flickr30k retrieval are reported with Recall@5. In Table˜28, we additionally report Recall@1 for all ViT-B/32 models from Table˜1. The overall ordering of methods is similar to Recall@5. Similarly, in Table˜27, we additionally report Recall@1 for all ViT-L/14 models from Table˜3.
Visual examples and failure case analysis To illustrate the strengths and limitations of the method and the compositionality task in general, we selected four of the SugarCrepe++ categories: Replace attribute, Replace relation, Swap attribute, and Swap object. For each, we examined the first instance in which the pretrained model’s prediction changed from incorrect to correct with CLIC-RedCaps, and vice versa. The results show that CLIC has lower absolute cosine similarity values compared to the pretrained model.
: A person holding up a chocolate doughnut with a face drawn on it.
: A chocolate doughnut with a face drawn on it is being held up by a person.
: A person holding up a vanilla doughnut with a face drawn on it.
Algorithm
CLIP (✗)
0.323
0.306
0.309
CLIC (✓)
0.280
0.285
0.263
: A tan toilet and sink combination in a small room.
: A small room with a tan toilet and sink combination positioned in close proximity to one another.
: A white toilet and sink combination in a small room.
Algorithm
CLIP (✓)
0.325
0.341
0.319
CLIC (✗)
0.260
0.256
0.262
: A pizza covered in lots of greens on top of a table.
: A table has a pizza on top of it covered in lots of greens.
: A pizza covered in lots of greens next to a table.
Algorithm
CLIP (✗)
0.300
0.280
0.293
CLIC (✓)
0.276
0.275
0.272
: A zebra is standing in an open field.
: On an open field, stands a zebra.
: A zebra is running across an open field.
Algorithm
CLIP (✓)
0.311
0.312
0.298
CLIC (✗)
0.283
0.254
0.270
: One apple and several oranges sit in a bowl.
: several oranges and an apple are positioned in a bowl.
: Several apples and one orange sit in a bowl.
Algorithm
CLIP (✗)
0.306
0.315
0.309
CLIC (✓)
0.276
0.298
0.270
: Two giraffe and a zebra are standing in a field.
: In a field, two giraffes and a zebra are standing.
: A giraffe and two zebras are standing in a field.
Algorithm
CLIP (✓)
0.316
0.316
0.315
CLIC (✗)
0.305
0.290
0.295
: Girls wash a motorcycle while men look on.
: The motorcycle is being washed by girls while men observe.
: Men wash a motorcycle while girls look on.
Algorithm
CLIP (✗)
0.365
0.356
0.358
CLIC (✓)
0.364
0.371
0.362
: A painting of a vase with a sunflower on a table.
: A vase containing a sunflower is positioned on a table in a painting.
: A painting of a sunflower with a vase on a table.
Algorithm
CLIP (✓)
0.378
0.381
0.378
CLIC (✗)
0.317
0.321
0.322